Differential equation question

please I need help with this question. Please show all work with clear hand writing

Find the solution of the second order differential equations: day + y = 0, y(TT/3) 0, y(TT/3) = 4 dx2 a. = b. y – 8y + 16y
Answer #1

u The given differential Equation is day dee? + y = 0 with initial condition y(1113) =0 f yenle) = 4 So, quvilalęy equationput cn = 2 in ez? O then 0+(2-0 becomes Ito (2) = 0 (15 - 26 so, finally put values of C1 = - 2 Jo ļ 62= 2 in y(a) then ya) =---------------------------------------------------/----------------------------------

so ito general Solution is 412) - (9+(20) Jee Here the given differential equation is y -8y +164 = 0 with initial condition

but we know Y10) -1, 50, put a zo in ytre) then y(x) =(1+62.2)e for be comes Y(x=0) = ((1+(2.0) x 0 = (atole y=0) = a Ş: 410)

Note: In example (b) we use x as independent variable